# Stable-RAG: Mitigating Retrieval-Permutation-Induced Hallucinations in Retrieval-Augmented Generation

### v1

@@ Abstract
Retrieval-Augmented Generation (RAG) has become a key paradigm for reducing factual hallucinations in large language models (LLMs), yet little is known about how the order of retrieved documents affects model behavior. We empirically show that under Top-5 retrieval with the gold document included, LLM answers vary substantially across permutations of the retrieved set, even when the gold document is fixed in the first position. This reveals a previously underexplored sensitivity to retrieval permutations. Although robust RAG methods primarily focus on enhancing LLM robustness to low-quality retrieval and mitigating positional bias to distribute attention fairly over long contexts, neither approach directly addresses permutation sensitivity. In this paper, we propose Stable-RAG, which exploits permutation sensitivity estimation to mitigate permutation-induced hallucinations. Stable-RAG runs the generator under multiple retrieval orders, clusters hidden states, and decodes from a cluster-center representation that captures the dominant reasoning pattern. It then uses these reasoning results to align hallucinated outputs toward the correct answer, encouraging the model to produce consistent and accurate predictions across document permutations. Experiments on three QA datasets show that Stable-RAG significantly improves answer accuracy, reasoning consistency and robust generalization across datasets, retrievers, and input lengths compared with baselines  [^1] Our code is available at https://github.com/zqc1023/Stable-RAG..
@@ 1 Introduction
Large language models (LLMs) have achieved remarkable performance on language understanding and generation tasks, but still often generate confident yet incorrect statements, known as factual hallucinations fan2024survey, especially in knowledge-intensive settings chen2022gere; huang-etal-2023-learning. Retrieval-Augmented Generation (RAG) gao2023retrieval; lewis2020retrieval reduces factual hallucinations by grounding model outputs in externally retrieved documents rather than relying only on parametric knowledge, improving factuality, interpretability, and updatability without additional retraining zhou2024trustworthiness.
Despite these benefits, RAG systems are far from hallucination-free. We identify a critical but overlooked vulnerability in existing RAG systems: a strong sensitivity to the order of retrieved documents. When the retrieved content remains exactly the same, including the gold document, merely reordering them can lead the model to follow entirely different reasoning paths and produce inconsistent answers, referred to as Permutation-Induced Hallucinations. As shown in Figure 1, we retrieve the Top-5 documents zhu2024information; xu2024recomp and place the gold document at different positions, LLM answers vary substantially across retrieval permutations. Even when the gold document is fixed first, models may still ignore it and produce answers that conflict with the evidence. This reveals a previously underexplored sensitivity to retrieval permutations, even for such short contexts shorter than one thousand tokens.
Existing work on RAG robustness mainly focus on retrieval quality and positional bias. The former enhances LLM robustness to low-quality retrieval via uncertainty estimation and adversarial training, such as noise injection raat; retrobust of weak-relevant documents. The latter alleviates attention bias toward specific positions in long contexts, promoting more balanced use of retrieved documents mspoe; wang-etal-2025-position. However, these approaches overlook a critical issue: permutation sensitivity is neither caused by weakly relevant documents, because the input documents are the same, nor confined to long-context reasoning tasks, since only the Top-5 documents fall within one thousand tokens.
@@ 2 Related Work
RAG mitigates factual hallucinations in LLMs for knowledge-intensive tasks by providing explicit evidence from external documents lewis2020retrieval; fan2024survey; chen2022gere. Prior work on improving the robustness of RAG systems has primarily focused on enhancing retrieval quality wang-etal-2025-richrag; xu2024recomp or strengthening the generator’s robustness. For instance, Selective-Context li2023compressing, EXIT exit, and AdaComp zhang2024adacomp apply noise filtering to boost generation accuracy; RetRobust retrobust and RAAT raat expose the model to retrieval noise or irrelevant documents during training, enhancing robustness. However, these methods generally assume a stable document order and do not systematically assess its impact on reasoning. Although ATM zhu-etal-2024-atm considers order perturbations, it does not explicitly model reasoning trajectories across permutations, and thus cannot ensure consistency.
Additionally, another line of research focuses positional bias in long-context scenarios. Most LLMs use relative positional encodings peysakhovich2023attention, such as RoPE su2024roformer or ALiBi press2021train, which introduce systematic biases: early tokens receive excessive attention due to attention sinks xiaoefficient; guattention, while long-range decay favors recent tokens. Prior work mitigates these issues by modifying positional encodings mspoe; chen-etal-2024-fortify; lin2024mixture, adjusting causal masks, reweighting attention or hidden states hsieh-etal-2024-found, or using Pos2Distill wang-etal-2025-position to distill knowledge from advantageous to less favorable positions to promote fair attention across tokens. However, these methods mainly target long contexts or large document sets and do not explicitly address reasoning inconsistencies induced by different permutations of the same retrieved document set.
@@ 3.2 Permutation Sensitivity Estimation
Recent work liang2025clue; lee-etal-2025-efficient-latent exploits hidden states to uncover latent reasoning trajectories, often as indicators of generative uncertainty. Accordingly, we propose to quantify model generation uncertainty via spectral clustering of hidden states. In this section, we validate the feasibility of spectral clustering algorithm through both layer-wise visualization and quantitative analysis.
@@ 4.1 Hidden State Clustering
For each query $q$ and its retrieved document set $\mathcal{S}=\{d_{1},d_{2},\dots,d_{n}\}$, we enumerate all permutations of the documents and run the model for each permutation. Let $i\in\{1,\dots,N\}$ denote the permutation index, where $N=n!$. To reduce computational cost, we extract only the last token hidden state of the final layer before response generation, $h^{(i)}\in\mathbb{R}^{d}$. Prior work azaria-mitchell-2023-internal; ni-etal-2025-towards has shown that this hidden state sufficiently captures the model’s perception of its knowledge boundaries. We organize all hidden states into a matrix $H$:
@@
To determine the number of clusters adaptively and capture the global structure of the hidden state space, we apply spectral clustering ng2001spectral to $H$, where each cluster corresponds to a latent reasoning mode lee-etal-2025-efficient-latent. We compute the similarity between each pair of hidden states $h^{(i)}$ and $h^{(j)}$ using the exponential of the cosine distance:
@@ 4.3 Alignment with DPO
We employ Direct Preference Optimization (DPO) rafailov2023direct to train the base model on the constructed preference tuples. For each tuple $(x,y_{w},y_{l})$, DPO maximizes the likelihood of the preferred answer $y_{w}$ over the less preferred $y_{l}$:
@@ 5.1 Experiments Setup
We evaluate our method on three QA benchmark datasets, including (1) Open-Domain QA, represented by NaturalQuestions (NQ) nq and TriviaQA triviaqa; (2) Multi-Hop QA, represented by HotpotQA hotpotqa. Dataset statistics are provided in Appendix A.1.
@@
Since answer style mismatch may cause additional variance, we follow prior work zhu-etal-2024-atm; peng-etal-2025-cafe; compselect and adopt Substring Exact Match (SubEM) and F1 for evaluation. SubEM checks whether the gold answer appears as a substring in the prediction, while F1 measures token-level overlap with the reference.
@@
We compare our method with the following baseline strategies on the same test set. Vanilla methods include Direct Generation, Vanilla RAG lewis2020retrieval, and Vanilla SFT zhang-etal-2024-r. Robust RAG methods include RetRobust retrobust, ATM zhu-etal-2024-atm, and RAAT raat. Positional Bias methods include Pos2Distill wang-etal-2025-position and Ms-PoE mspoe. The details of these baselines are presented in Appendix A.2.
@@
We use LLaMA3-8B-Instruct llama3 and Qwen3-8B qwen3 as backbone models for experiments. To ensure high and consistent evaluation quality cuconasu2024power and further assess the stability of our method under different retrieval settings, we follow prior work zhu2024information; xu2024recomp; refiner and use the same Top-5 Wikipedia passages retrieved by DPR dpr and Contriever-MS MARCO izacard2021unsupervised for all baselines and our method. Additional implementation details are provided in Appendix A.3.
@@ 6 Conclusion
Acknowledgments
This work was funded by the National Natural Science Foundation of China (NSFC) under Grants No. 62406013, the Beijing Advanced Innovation Center Funds for Future Blockchain and Privacy Computing(GJJ-24-034), and the Fundamental Research Funds for the Central Universities.
@@ A.1 Datasets
We conduct experiments on three widely used QA datasets that cover both single-hop and multi-hop question-answering scenarios. Table 6 summarizes the key statistics of these datasets. Specifically, NQ nq and TriviaQA triviaqa are representative single-hop datasets, where each question can typically be answered using information from a single passage retrieved from the corpus. These datasets primarily evaluate a model’s ability to locate and extract factual evidence efficiently. In contrast, HotpotQA hotpotqa is a challenging multi-hop dataset that requires integrating and reasoning over multiple pieces of evidence distributed across different documents to derive the final answer. This dataset is particularly useful for testing a model’s reasoning and compositional understanding capabilities. Together, these datasets provide a comprehensive benchmark for evaluating both the retrieval quality and reasoning robustness of our proposed method under diverse task settings.
@@ A.2 Baseline Details
(i) Direct Generation. This baseline relies solely on the generator’s parametric knowledge to produce answers without consulting any retrieved documents. (ii) Vanilla RAG lewis2020retrieval. This baseline concatenates all retrieved documents as model input without any additional processing. (iii) Vanilla SFT. We implement vanillla SFT following  zhang-etal-2024-r. For each training example, this baseline uses the gold answer as the training label if it appears in the retrieved documents; otherwise, it assigns “I don’t know” as the training label to guide the model to abstain when the necessary information is missing.
@@
(i) RetRobust retrobust. This baseline improves retrieval-augmented QA models by filtering out irrelevant retrieved passages and fine-tuning the model on a mix of relevant and irrelevant contexts, enabling it to leverage relevant information while remaining robust to irrelevant content. (ii) ATM zhu-etal-2024-atm. This baseline optimizes a retrieval-augmented Generator using an Adversarial Tuning Multi-agent system, where an auxiliary Attacker agent iteratively steers the Generator to better discriminate useful documents from noisy or fabricated ones, improving robustness and performance on knowledge-intensive question answering tasks. (iii) RAAT raat. This baseline dynamically adjusts the model’s learning process in response to various types of retrieval noise through adaptive adversarial training, while employing multi-task learning to enable the model to internally recognize and handle noisy contexts, thereby improving robustness and answer quality in retrieval-augmented generation.
@@
(i) Pos2Distill wang-etal-2025-position. This baseline mitigates positional bias in long-context tasks by transferring knowledge from advantageous positions to less favorable ones through position-to-position knowledge distillation. (ii) Ms-PoE mspoe. This baseline uses Multi-scale Positional Encoding to mitigate the "lost-in-the-middle" issue in LLMs by rescaling positional indices and assigning different scaling ratios to attention heads, enabling multi-scale context fusion without fine-tuning or extra overhead.
@@ A.3 Training Details
We use LLaMA3-8B-Instruct  [^2] https://huggingface.co/meta-llama/Meta-Llama-3-8B-Instruct llama3 and Qwen3-8B  [^3] https://huggingface.co/Qwen/Qwen3-8B qwen3 as backbone models for experiments. We implement our DPO training pipeline using the HuggingFace Transformers wolf-etal-2020-transformers and TRL libraries vonwerra2022trl, incorporating PEFT LoRA hu2022lora for parameter-efficient fine-tuning. Both the base model and reference model are initialized from pre-trained checkpoints, with the reference model kept in evaluation mode to provide stable policy targets during training. Each dataset is randomly shuffled and split into 85% training and 15% validation samples, with a maximum of 18,000 samples per dataset to control computational overhead. To guarantee reproducible results, we use a fixed random seed with a value of 42. LoRA is applied to all projection layers, including query, key, value, output, gate, up, and down projections, with rank $r=128$, alpha =128=128, dropout =0=0 and no additional bias terms. The DPO configuration uses a per-device batch size of 2 with gradient accumulation of 8, a learning rate of $5\times 10^{-6}$, a linear warmup ratio of 0.1, and a preference scaling hyperparameter $\beta$ of 0.4. We train LLaMA-3-8B-Instruct for a single epoch and Qwen3-8B for two epochs on two NVIDIA RTX PRO 6000 GPUs, with each epoch taking roughly two hours. After training, the fine-tuned models and tokenizers are saved for downstream evaluation.
@@ B.1 Spectral Clustering on Hidden States
to adaptively determine the number of clusters and capture the global structure of the hidden state space, where each cluster corresponds to a latent reasoning mode lee-etal-2025-efficient-latent.
@@ B.4 Eigen-decomposition and Determining Cluster Number
ensuring clear separation between latent reasoning modes following tandard practice ng2001spectral; von2007tutorial.

### v2

@@ Abstract
Retrieval-Augmented Generation (RAG) has become a key paradigm for reducing factual hallucinations in large language models (LLMs), yet little is known about how the order of retrieved documents affects model behavior. We empirically show that under Top-5 retrieval with the gold document included, LLM answers vary substantially across permutations of the retrieved set, even when the gold document is fixed in the first position. This reveals a previously underexplored sensitivity to retrieval permutations. Although robust RAG methods primarily focus on enhancing LLM robustness to low-quality retrieval and mitigating positional bias to distribute attention fairly over long contexts, neither approach directly addresses permutation sensitivity. In this paper, we propose Stable-RAG, which exploits permutation sensitivity estimation to mitigate permutation-induced hallucinations. Stable-RAG runs the generator under multiple retrieval orders, clusters hidden states, and decodes from a cluster-center representation that captures the dominant reasoning pattern. It then uses these reasoning results to align hallucinated outputs toward the correct answer, encouraging the model to produce consistent and accurate predictions across document permutations. Experiments on three QA datasets show that Stable-RAG significantly improves answer accuracy, reasoning consistency and robust generalization across datasets, retrievers, and input lengths compared with baselines.
@@ 1 Introduction
Large language models (LLMs) have achieved remarkable performance on language understanding and generation tasks, but still often generate confident yet incorrect statements, known as factual hallucinations Fan et al. (2024), especially in knowledge-intensive settings Chen et al. (2022); Huang et al. (2023). Retrieval-Augmented Generation (RAG) Gao et al. (2023); Lewis et al. (2020) reduces factual hallucinations by grounding model outputs in externally retrieved documents rather than relying only on parametric knowledge, improving factuality, interpretability, and updatability without additional retraining Zhou et al. (2024).
Despite these benefits, RAG systems are far from hallucination-free. We identify a critical but overlooked vulnerability in existing RAG systems: a strong sensitivity to the order of retrieved documents. When the retrieved content remains exactly the same, including the gold document, merely reordering them can lead the model to follow entirely different reasoning paths and produce inconsistent answers, referred to as Permutation-Induced Hallucinations. As shown in Figure 1, we retrieve the Top-5 documents Zhu et al. (2024b); Xu et al. (2024) and place the gold document at different positions, LLM answers vary substantially across retrieval permutations. Even when the gold document is fixed first, models may still ignore it and produce answers that conflict with the evidence. This reveals a previously underexplored sensitivity to retrieval permutations, even for such short contexts shorter than one thousand tokens.
Existing work on RAG robustness mainly focus on retrieval quality and positional bias. The former enhances LLM robustness to low-quality retrieval via uncertainty estimation and adversarial training, such as noise injection Fang et al. (2024); Yoran et al. (2024) of weak-relevant documents. The latter alleviates attention bias toward specific positions in long contexts, promoting more balanced use of retrieved documents Zhang et al. (2024c); Wang et al. (2025b). However, these approaches overlook a critical issue: permutation sensitivity is neither caused by weakly relevant documents, because the input documents are the same, nor confined to long-context reasoning tasks, since only the Top-5 documents fall within one thousand tokens.
@@ 2 Related Work
RAG mitigates factual hallucinations in LLMs for knowledge-intensive tasks by providing explicit evidence from external documents Lewis et al. (2020); Fan et al. (2024); Chen et al. (2022). Prior work on improving the robustness of RAG systems has primarily focused on enhancing retrieval quality Wang et al. (2025a); Xu et al. (2024) or strengthening the generator’s robustness. For instance, Selective-Context Li et al. (2023), EXIT Hwang et al. (2025), and AdaComp Zhang et al. (2024b) apply noise filtering to boost generation accuracy; RetRobust Yoran et al. (2024) and RAAT Fang et al. (2024) expose the model to retrieval noise or irrelevant documents during training, enhancing robustness. However, these methods generally assume a stable document order and do not systematically assess its impact on reasoning. Although ATM Zhu et al. (2024a) considers order perturbations, it does not explicitly model reasoning trajectories across permutations, and thus cannot ensure consistency.
Additionally, another line of research focuses positional bias in long-context scenarios. Most LLMs use relative positional encodings Peysakhovich and Lerer (2023), such as RoPE Su et al. (2024) or ALiBi Press et al. (2021), which introduce systematic biases: early tokens receive excessive attention due to attention sinks Xiao et al. ; Gu et al. , while long-range decay favors recent tokens. Prior work mitigates these issues by modifying positional encodings Zhang et al. (2024c); Chen et al. (2024); Lin et al. (2024), adjusting causal masks, reweighting attention or hidden states Hsieh et al. (2024), or using Pos2Distill Wang et al. (2025b) to distill knowledge from advantageous to less favorable positions to promote fair attention across tokens. However, these methods mainly target long contexts or large document sets and do not explicitly address reasoning inconsistencies induced by different permutations of the same retrieved document set.
@@ 3.2 Permutation Sensitivity Estimation
Recent work Liang et al. (2025); Lee et al. (2025) exploits hidden states to uncover latent reasoning trajectories, often as indicators of generative uncertainty. Accordingly, we propose to quantify model generation uncertainty via spectral clustering of hidden states. In this section, we validate the feasibility of spectral clustering algorithm through both layer-wise visualization and quantitative analysis.
@@ 4.1 Hidden State Clustering
For each query $q$ and its retrieved document set $\mathcal{S}=\{d_{1},d_{2},\dots,d_{n}\}$, we enumerate all permutations of the documents and run the model for each permutation. Let $i\in\{1,\dots,N\}$ denote the permutation index, where $N=n!$. To reduce computational cost, we extract only the last token hidden state of the final layer before response generation, $h^{(i)}\in\mathbb{R}^{d}$. Prior work Azaria and Mitchell (2023); Ni et al. (2025) has shown that this hidden state sufficiently captures the model’s perception of its knowledge boundaries. We organize all hidden states into a matrix $H$:
@@
To determine the number of clusters adaptively and capture the global structure of the hidden state space, we apply spectral clustering Ng et al. (2001) to $H$, where each cluster corresponds to a latent reasoning mode Lee et al. (2025). We compute the similarity between each pair of hidden states $h^{(i)}$ and $h^{(j)}$ using the exponential of the cosine distance:
@@ 4.3 Alignment with DPO
We employ Direct Preference Optimization (DPO) Rafailov et al. (2023) to train the base model on the constructed preference tuples. For each tuple $(x,y_{w},y_{l})$, DPO maximizes the likelihood of the preferred answer $y_{w}$ over the less preferred $y_{l}$:
@@ 5.1 Experiments Setup
We evaluate our method on three QA benchmark datasets, including (1) Open-Domain QA, represented by NaturalQuestions (NQ) Kwiatkowski et al. (2019) and TriviaQA Joshi et al. (2017); (2) Multi-Hop QA, represented by HotpotQA Yang et al. (2018). Dataset statistics are provided in Appendix A.1.
@@
Since answer style mismatch may cause additional variance, we follow prior work Zhu et al. (2024a); Peng et al. (2025); Zhang et al. (2025) and adopt Substring Exact Match (SubEM) and F1 for evaluation. SubEM checks whether the gold answer appears as a substring in the prediction, while F1 measures token-level overlap with the reference.
@@
We compare our method with the following baseline strategies on the same test set. Vanilla methods include Direct Generation, Vanilla RAG Lewis et al. (2020), and Vanilla SFT Zhang et al. (2024a). Robust RAG methods include RetRobust Yoran et al. (2024), ATM Zhu et al. (2024a), and RAAT Fang et al. (2024). Positional Bias methods include Pos2Distill Wang et al. (2025b) and Ms-PoE Zhang et al. (2024c). The details of these baselines are presented in Appendix A.2.
@@
We use LLaMA3-8B-Instruct Dubey et al. (2024) and Qwen3-8B Yang et al. (2025) as backbone models for experiments. To ensure high and consistent evaluation quality Cuconasu et al. (2024) and further assess the stability of our method under different retrieval settings, we follow prior work Zhu et al. (2024b); Xu et al. (2024); Li et al. (2024) and use the same Top-5 Wikipedia passages retrieved by DPR Karpukhin et al. (2020) and Contriever-MS MARCO Izacard et al. (2021) for all baselines and our method. Additional implementation details are provided in Appendix A.3.
@@ A.1 Datasets
We conduct experiments on three widely used QA datasets that cover both single-hop and multi-hop question-answering scenarios. Table 6 summarizes the key statistics of these datasets. Specifically, NQ Kwiatkowski et al. (2019) and TriviaQA Joshi et al. (2017) are representative single-hop datasets, where each question can typically be answered using information from a single passage retrieved from the corpus. These datasets primarily evaluate a model’s ability to locate and extract factual evidence efficiently. In contrast, HotpotQA Yang et al. (2018) is a challenging multi-hop dataset that requires integrating and reasoning over multiple pieces of evidence distributed across different documents to derive the final answer. This dataset is particularly useful for testing a model’s reasoning and compositional understanding capabilities. Together, these datasets provide a comprehensive benchmark for evaluating both the retrieval quality and reasoning robustness of our proposed method under diverse task settings.
@@ A.2 Baseline Details
(i) Direct Generation. This baseline relies solely on the generator’s parametric knowledge to produce answers without consulting any retrieved documents. (ii) Vanilla RAG Lewis et al. (2020). This baseline concatenates all retrieved documents as model input without any additional processing. (iii) Vanilla SFT. We implement vanillla SFT following  Zhang et al. (2024a). For each training example, this baseline uses the gold answer as the training label if it appears in the retrieved documents; otherwise, it assigns “I don’t know” as the training label to guide the model to abstain when the necessary information is missing.
@@
(i) RetRobust Yoran et al. (2024). This baseline improves retrieval-augmented QA models by filtering out irrelevant retrieved passages and fine-tuning the model on a mix of relevant and irrelevant contexts, enabling it to leverage relevant information while remaining robust to irrelevant content. (ii) ATM Zhu et al. (2024a). This baseline optimizes a retrieval-augmented Generator using an Adversarial Tuning Multi-agent system, where an auxiliary Attacker agent iteratively steers the Generator to better discriminate useful documents from noisy or fabricated ones, improving robustness and performance on knowledge-intensive question answering tasks. (iii) RAAT Fang et al. (2024). This baseline dynamically adjusts the model’s learning process in response to various types of retrieval noise through adaptive adversarial training, while employing multi-task learning to enable the model to internally recognize and handle noisy contexts, thereby improving robustness and answer quality in retrieval-augmented generation.
@@
(i) Pos2Distill Wang et al. (2025b). This baseline mitigates positional bias in long-context tasks by transferring knowledge from advantageous positions to less favorable ones through position-to-position knowledge distillation. (ii) Ms-PoE Zhang et al. (2024c). This baseline uses Multi-scale Positional Encoding to mitigate the "lost-in-the-middle" issue in LLMs by rescaling positional indices and assigning different scaling ratios to attention heads, enabling multi-scale context fusion without fine-tuning or extra overhead.
@@ A.3 Training Details
We use LLaMA3-8B-Instruct  [^1] https://huggingface.co/meta-llama/Meta-Llama-3-8B-Instruct Dubey et al. (2024) and Qwen3-8B  [^2] https://huggingface.co/Qwen/Qwen3-8B Yang et al. (2025) as backbone models for experiments. We implement our DPO training pipeline using the HuggingFace Transformers Wolf et al. (2020) and TRL libraries von Werra et al. (2020), incorporating PEFT LoRA Hu et al. (2022) for parameter-efficient fine-tuning. Both the base model and reference model are initialized from pre-trained checkpoints, with the reference model kept in evaluation mode to provide stable policy targets during training. Each dataset is randomly shuffled and split into 85% training and 15% validation samples, with a maximum of 18,000 samples per dataset to control computational overhead. To guarantee reproducible results, we use a fixed random seed with a value of 42. LoRA is applied to all projection layers, including query, key, value, output, gate, up, and down projections, with rank $r=128$, alpha =128=128, dropout =0=0 and no additional bias terms. The DPO configuration uses a per-device batch size of 2 with gradient accumulation of 8, a learning rate of $5\times 10^{-6}$, a linear warmup ratio of 0.1, and a preference scaling hyperparameter $\beta$ of 0.4. We train LLaMA-3-8B-Instruct for a single epoch and Qwen3-8B for two epochs on two NVIDIA RTX PRO 6000 GPUs, with each epoch taking roughly two hours. After training, the fine-tuned models and tokenizers are saved for downstream evaluation.
@@ B.1 Spectral Clustering on Hidden States
to adaptively determine the number of clusters and capture the global structure of the hidden state space, where each cluster corresponds to a latent reasoning mode Lee et al. (2025).
@@ B.4 Eigen-decomposition and Determining Cluster Number
ensuring clear separation between latent reasoning modes following standard practice Ng et al. (2001); Von Luxburg (2007).